# Dirac quantum criticality in twisted double bilayer transition metal dichalcogenides

### v1

@@ VI Pressure-tuned phase diagram
The critical pressure depends sensitively on the twist angle. For example, if the offset from the ambient-pressure phase boundary doubles to $\theta-\theta_{\mathrm{c}}(0)=0.1^{\circ}$, the required pressure increases to $p_{\mathrm{c}}\simeq 0.6\,\mathrm{GPa}$, see Appendix C. This strong dependence highlights uniaxial pressure as a highly effective tuning knob for accessing and traversing the critical region above the Dirac-semimetal-to-antiferromagnetic-insulator quantum critical point.
@@ VIII.1 Strain-free samples
The Hartree-Fock analysis in Secs. V and VI revealed a continuous quantum phase transition between the symmetric Dirac semimetal and the antiferromagnetic insulator, driven by twist angle $\theta$ or uniaxial pressure $p$. At the quantum critical point, both the correlation length and correlation time diverge, so that the low-energy physics becomes universal and is fully captured by an effective continuum quantum field theory. This quantum field theory can be deduced from symmetry considerations: Define the eight-component Dirac fermion field $\psi_{\mathbf{k}}=(c_{\boldsymbol{\kappa}+\mathbf{k},n,s},c_{\boldsymbol{\kappa}^{\prime}+\mathbf{k},n,s})$, at long wavelengths $\mathbf{k}\ll 1/a_{\mathrm{M}}$, where $a_{\mathrm{M}}\sim a_{0}/\theta$ denotes the moiré lattice scale and $c_{\boldsymbol{\kappa}+\mathbf{k},n,s}$ ($c_{\boldsymbol{\kappa}’+\mathbf{k},n,s}$) annihilates an electron at wavevector $\mathbf{k}$ relative to the moiré valley $\boldsymbol{\kappa}$ ($\boldsymbol{\kappa}$’) with band index $n=1,2$ and spin $s={\uparrow,\downarrow}$. We introduce the antiferromagnetic order parameter $\vec{\varphi}$, constructed from the long-wavelength staggered magnetization. Under SU(2) spin rotations, $\vec{\varphi}$ transforms as a vector. Consequently, at low energies and long wavelengths, the system is described by the Gross-Neveu-Heisenberg action in $(2+1)$-dimensional Euclidean space-time [75],
@@
$S_{\text{GNH}}=\int\hskip-2.5pt\mathrm{d}\tau\mathrm{d}^{2}\vec{x}\,\biggl{[}\bar{\psi}(\gamma_{\mu}\otimes\mathbb{1}_{2})\partial_{\mu}\psi+g\vec{\varphi}\cdot\bar{\psi}(\mathbb{1}_{4}\otimes\vec{s})\psi\\ +\frac{1}{2}(\partial_{\mu}\vec{\varphi})^{2}+\frac{r}{2}\vec{\varphi}^{2}+\lambda\bigl{(}\vec{\varphi}^{2}\bigr{)}^{2}\biggr{]}\,$, (30)
with summation over $\mu=0,1,2$, where $(\partial_{\mu})=(\partial_{\tau},\partial_{x},\partial_{y})$ is the $(2+1)$-dimensional space-time derivative. The $\gamma_{\mu}$ are $4\times 4$ Dirac matrices acting on band and valley indices, satisfying $\{\gamma_{\mu},\gamma_{\nu}\}=2\delta_{\mu\nu}$, and $\vec{s}=(s_{x},s_{y},s_{z})$ are Pauli matrices acting on spin. The tuning parameter $r$ varies linearly with $\theta-\theta_{\mathrm{c}}$ or $p_{\mathrm{c}}-p$ near the transition, while $g$ and $\lambda$ denote the Yukawa coupling and boson self-interaction, respectively. The Gross-Neveu-Heisenberg model defined in Eq. (30) possesses explicit SU(2) spin symmetry and emergent relativistic invariance in $(2+1)$-dimensional space-time [76].
@@
For $r>0$, the mean-field solution yields an SU(2)-symmetric phase with vanishing order parameter $\langle\vec{\varphi}\rangle=0$ and an electronic spectrum hosting two gapless Dirac cones per spin. For $r<0$, the order parameter develops a finite expectation value $\langle\vec{\varphi}\rangle\neq 0$, gapping out the Dirac cones and giving rise to a phase with long-range antiferromagnetic order.
Beyond mean-field theory, the critical behavior of the Gross-Neveu-Heisenberg model has been investigated using various analytical approaches, including the $\epsilon$ expansion around the upper [3, 77] and lower [78] critical dimensions, the large-$N$ expansion [79], and the functional renormalization group [5, 80]. Complementary numerical studies have investigated the transition using different lattice realizations [4, 81, 82, 83, 6, 84, 85, 86, 87, 88, 89, 90]. The thermodynamic, spectroscopic, and transport properties near the quantum critical point are governed by a set of universal critical exponents. In the antiferromagnetic phase, the staggered magnetization scales with twist angle or pressure as $\langle\vec{\varphi}\rangle\propto(\theta_{\mathrm{c}}-\theta)^{\beta}$ or $\langle\vec{\varphi}\rangle\propto(p-p_{\mathrm{c}})^{\beta}$, with nonuniversal prefactors but a universal order-parameter exponent $\beta\approx 1.21$. The electronic gap exhibits the scaling $\Delta\propto(\theta_{\mathrm{c}}-\theta)^{\nu z}$ or $\Delta\propto(p-p_{\mathrm{c}})^{\nu z}$, with dynamical exponent $z=1$ due to the emergent relativistic symmetry and correlation-length exponent $\nu\approx 1.20$. At the quantum critical point $\theta=\theta_{\mathrm{c}}$ or $p=p_{\mathrm{c}}$, the dynamical spin structure factor scales as $\mathcal{S}(\mathbf{k},\omega)\propto 1/{(\omega^{2}-\mathbf{k}^{2})^{(2-\eta_{\varphi})/2}}$, with order-parameter anomalous dimension $\eta_{\varphi}\approx 1.01$. The quoted exponents are estimated from interpolations between expansions near the lower and upper critical dimensions [78], using standard hyperscaling relations [91].
@@ VIII.2 Strained samples
We emphasize that the O(3) model above is valid only at energies below the strain-induced single-particle gap $\Delta_{\epsilon}$. For weakly strained samples, $\Delta_{\epsilon}$ is small compared to the microscopic interaction scale $J$, which sets the order of the interaction-induced band gap $\Delta$ deep in the ordered phase. In the strong-coupling regime, this scale can be estimated as $J\sim D^{2}/V$, with $V\sim e^{2}/(\epsilon_{0}\epsilon_{\text{eff}}a_{\mathrm{M}})$ the effective interaction strength, $a_{\mathrm{M}}\sim a_{0}/\theta$ the moiré length scale, and $D$ the low-energy bandwidth. For twist angles near the critical value, $\theta\sim\theta_{\mathrm{c}}=2.7^{\circ}$, we estimate $D\sim 20\,\text{meV}$ and $V\sim 20\,\text{meV}$ for $\epsilon_{\text{eff}}\sim 100$ and $a_{\mathrm{M}}\sim 10\,\text{nm}$, giving $J\sim 20\,\text{meV}$ for the microscopic energy scale. At energies above $\Delta_{\epsilon}$ but below $J$, such systems remain governed by Gross-Neveu-Heisenberg universality, described by Eq. (30). This leads to a crossover within the quantum critical regime at finite temperatures $T$ above the quantum critical point at $\theta=\theta_{\mathrm{c}}$ or $p=p_{\mathrm{c}}$, as schematically shown in Fig. 7.

### v2

@@ VI Pressure-tuned phase diagram
The critical pressure depends sensitively on the twist angle. For example, if the offset from the ambient-pressure phase boundary doubles to $\theta-\theta_{\mathrm{c}}(0)=0.1^{\circ}$, the required pressure increases to $p_{\mathrm{c}}\simeq 0.6\,\mathrm{GPa}$, see Appendix D. This strong dependence highlights uniaxial pressure as a highly effective tuning knob for accessing and traversing the critical region above the Dirac-semimetal-to-antiferromagnetic-insulator quantum critical point.
@@ VIII.1 Strain-free samples
The Hartree-Fock analysis in Secs. V and VI suggested a continuous quantum phase transition between the symmetric Dirac semimetal and the antiferromagnetic insulator, driven by twist angle $\theta$ or uniaxial pressure $p$. Near the phase transition, however, fluctuations become important and Hartree-Fock theory may no longer be adequate. To accurately describe this physics, we derive the most general effective low-energy field theory consistent with the microscopic symmetries of the model and the nature of the adjacent phases, suitable for further analysis within a renormalization group approach. This quantum field theory can be deduced from symmetry considerations: Define the eight-component Dirac fermion field $\psi_{\mathbf{k}}=(c_{\boldsymbol{\kappa}+\mathbf{k},n,s},c_{\boldsymbol{\kappa}^{\prime}+\mathbf{k},n,s})$, at long wavelengths $\mathbf{k}\ll\pi/a_{\mathrm{M}}$, where $a_{\mathrm{M}}\simeq a_{0}/\theta$ denotes the moiré lattice scale and $c_{\boldsymbol{\kappa}+\mathbf{k},n,s}$ ($c_{\boldsymbol{\kappa}^{\prime}+\mathbf{k},n,s}$) annihilates an electron at wavevector $\mathbf{k}$ relative to the moiré minivalley $\boldsymbol{\kappa}$ ($\boldsymbol{\kappa}$’) with band index $n=1,2$ and spin $s={\uparrow,\downarrow}$. We introduce the antiferromagnetic order parameter $\vec{\varphi}$, constructed from the long-wavelength staggered magnetization. Under SU(2) spin rotations, $\vec{\varphi}$ transforms as a vector. Consequently, the most general symmetry-allowed low-energy effective Lagrangian takes the form
@@
$\mathcal{L}=\psi^{\dagger}(\partial_{\tau}+v_{\mathrm{F}}\gamma_{0}\vec{\gamma}\cdot\nabla)\psi+\frac{1}{2}\vec{\varphi}\cdot(r-\partial_{\tau}^{2}-v_{\mathrm{B}}^{2}\Delta)\vec{\varphi}\\ +g\vec{\varphi}\cdot\psi^{\dagger}\gamma_{0}\vec{s}\psi+\lambda\bigl(\vec{\varphi}\cdot\vec{\varphi}\bigr)^{2}+\mathcal{O}(\varphi^{6},\psi^{4},\psi^{2}\varphi^{4})\,$, (30)
where $\psi\equiv\psi(\vec{x},\tau)=k_{\mathrm{B}}T\sum_{n}\int^{\Lambda}_{0}\hskip-2.5pt\frac{\mathrm{d}^{2}\vec{k}}{(2\pi a_{\mathrm{M}})^{2}}\mathrm{e}^{\mathrm{i}\omega_{n}\tau+i\mathbf{k}\cdot\vec{x}}\psi_{\mathbf{k}}(\omega_{n})$, with fermionic Matsubara frequencies $\hbar\omega_{n}=k_{\mathrm{B}}T(2n+1)\pi$ at temperature $T$ and an ultraviolet momentum cutoff $\Lambda\simeq 1/a_{\mathrm{M}}$. The velocities $v_{\mathrm{F}}$ and $v_{\mathrm{B}}$ characterize the propagation speeds of the low-energy fermionic and magnetic excitations, respectively. The tuning parameter $r$ varies linearly with $\theta-\theta_{\mathrm{c}}$ or $p_{\mathrm{c}}-p$ in the vicinity of the transition, while $g$ and $\lambda$ denote the Yukawa coupling and boson self-interaction, respectively. The higher-order interaction terms indicated in Eq. (30) are irrelevant in the renormalization-group sense and therefore do not influence the universal critical behavior. The $\gamma_{\mu}$ are $4\times 4$ Dirac matrices acting on minivalley and sublattice degrees of freedom and satisfy the Clifford algebra $\{\gamma_{\mu},\gamma_{\nu}\}=2\delta_{\mu\nu}$. In the reference frame defined by $\mathbf{k}_{x}=\mathbf{k}\cdot\boldsymbol{\kappa}/\kappa$ and $k_{y}=(\boldsymbol{\kappa}\times\mathbf{k})\times\boldsymbol{\kappa}/\kappa^{2}$, a convenient representation of the Dirac matrices is $(\gamma_{0},\gamma_{1},\gamma_{2})=(\tau_{0}\sigma_{z},\tau_{z}\sigma_{y},\tau_{0}\sigma_{x})$ [2]. Here, $(\tau_{x},\tau_{y},\tau_{z})$ denote Pauli matrices in minivalley space with $\tau_{0}$ the identity, while $(\sigma_{x},\sigma_{y},\sigma_{z})$ act in sublattice space. $\vec{s}=(s_{x},s_{y},s_{z})$ denotes the Pauli matrices in spin space. For generic values of $v_{\mathrm{F}}\neq v_{\mathrm{B}}$, the Lagrangian lacks Lorentz symmetry. At the phase transition point, however, we show below that $v_{\mathrm{F}}$ and $v_{\mathrm{B}}$ flow to the same value at low energies, giving rise to emergent Lorentz invariance at criticality.
@@
For $r>0$, the mean-field solution to the theory defined in Eq. (1) describes an SU(2)-symmetric phase with a vanishing order parameter, $\langle\vec{\varphi}\rangle=0$, and an electronic spectrum featuring two gapless Dirac cones per spin. For $r<0$, the order parameter acquires a finite expectation value at zero temperature, $\langle\vec{\varphi}\rangle\neq 0$, which gaps out the Dirac cones and yields a phase with long-range antiferromagnetic order.
To describe the physics near the phase transition beyond mean-field theory, we employ a renormalization group analysis based on an $\epsilon$ expansion about the upper critical spatial dimension $d=3$ [78]. As a starting point of this analysis, we estimate the values of the theory parameters at the moiré lattice scale $\Lambda\simeq 1/a_{\mathrm{M}}$. From the spectrum of $\mathcal{H}_{\text{kin}}$ in Eq. (1), we extract a Fermi velocity $v_{\mathrm{F}}\simeq 2.7\times 10^{4}\,\mathrm{m/s}$ for $\theta\simeq 2.7^{\circ}$, in agreement to within 10% with the value inferred from magnetotransport measurements in twisted double bilayer WSe2 for $\theta\simeq 3^{\circ}$ [35]. To estimate $v_{\mathrm{B}}$, we consider an effective Hubbard model on the emergent honeycomb lattice in the strong-coupling limit, which yields an antiferromagnetic Heisenberg exchange coupling $J\simeq 4t_{1}^{2}/U_{0}$ [79], where $t_{1}$ is the effective nearest-neighbor hopping amplitude and $U_{0}$ is the effective on-site interaction. The magnetic excitation spectrum contains a gapless magnon mode with linear dispersion $\omega_{\mathbf{k}}=v_{\mathrm{B}}|\mathbf{k}|+\mathcal{O}(k^{2})$ near the Brillouin-zone center. Within linear spin-wave theory, the corresponding magnon velocity is $v_{\mathrm{B}}\simeq\sqrt{3/2}a_{\text{M}}JS/\hbar$ [80], where $a_{\text{M}}\simeq a_{0}/\theta$ is the moiré lattice constant and $S=1/2$ is the spin length. Within the effective Hubbard-model description, the Fermi velocity in the weak-coupling limit is $v_{\mathrm{F}}=\sqrt{3}t_{1}a_{\mathrm{M}}/(2\hbar)$, which allows for a rough order-of-magnitude estimate of the spin-wave velocity relative to the Fermi velocity as $v_{\mathrm{B}}\simeq 4\sqrt{2}(t_{1}/U_{0})Sv_{\mathrm{F}}\simeq 2\times 10^{4}\,\mathrm{m/s}$. In the last step, we have used the critical ratio $(U_{0}/t_{1})_{\mathrm{c}}\simeq 3.8$ for the semimetal-to-insulator transition of the honeycomb-lattice Hubbard model [4], together with the numerical value of $v_{\mathrm{F}}$ extracted from the spectrum of $\mathcal{H}_{\text{kin}}$. The ratio $g^{2}/(2r)$ sets the effective electron-electron interaction scale responsible for the antiferromagnetic order. In Appendix B.3, we show that for $\theta=2.7^{\circ}$ and $\epsilon_{\text{eff}}=110$ this scale can be estimated as $g^{2}/(2r)\simeq U_{\mathrm{S}}\simeq 4.6\,\mathrm{meV}$. The bosonic self-coupling $\lambda$ can be taken to be small at the moiré lattice scale; however, it is generated under renormalization group flow and must therefore be included to correctly capture the low-energy physics.
Integrating out modes in the momentum shell $\mathbf{k}\in[\Lambda/b,\Lambda]$ in $d=3-\varepsilon$ spatial dimensions, together with all Matsubara frequencies $\omega\in(-\infty,\infty)$ at zero temperature, induces a renormalization group flow of the parameters in Eq. (1) along the critical surface $r=r_{\star}$, given by
$\frac{\mathrm{d}(v_{\mathrm{F}}/v_{\mathrm{B}})}{\mathrm{d}\ln b}$ $=(z-1-\eta_{\psi})v_{\mathrm{F}}/v_{\mathrm{B}}+\frac{3}{8}f_{p}(v_{\mathrm{F}}/v_{\mathrm{B}})g^{2}\,$, (31)
$\frac{\mathrm{d}g^{2}}{\mathrm{d}\ln b}$ $=(\varepsilon+z-1-\eta_{1}-2\eta_{\psi})g^{2}+\frac{1}{2}f_{g}(v_{\mathrm{F}}/v_{\mathrm{B}})g^{4}\,$, (32)
$\frac{\mathrm{d}\lambda}{\mathrm{d}\ln b}$ $=(\varepsilon+3z-3-2\eta_{\phi})\lambda-11\lambda^{2}+\frac{1}{2}f_{\lambda}(v_{\mathrm{F}}/v_{\mathrm{B}})g^{2}\,$, (33)
where the dynamical critical exponent $z$ and the anomalous dimensions are defined as
$z$ $=1+\frac{\eta_{\phi}-\eta_{1}}{2}\,$, $\eta_{\psi}$ $=\frac{3}{8}f_{\omega}(v_{\mathrm{F}}/v_{\mathrm{B}})g^{2}\,$, (34)
$\eta_{\phi}$ $=f_{\omega^{2}}(v_{\mathrm{F}}/v_{\mathrm{B}})g^{2}\,$, $\eta_{1}$ $=f_{p^{2}}(v_{\mathrm{F}}/v_{\mathrm{B}})g^{2}\,.$ (35)
Away from the critical surface, the tuning parameter $r$ evolves as
$\frac{\mathrm{d}r}{\mathrm{d}\ln b}$ $=(2-\eta_{1})r+10\frac{\lambda}{\sqrt{1+r}}-4f_{r}(v_{\mathrm{F}}/v_{\mathrm{B}})g^{2}\,.$ (36)
In the above equations, we have introduced dimensionless parameters according to $g^{2}\Lambda^{1-z+\eta_{1}+2\eta_{\psi}-\varepsilon}/(2\pi^{2})\mapsto g^{2}$, $\lambda\Lambda^{3-3z+2\eta_{\phi}-\varepsilon}/(2\pi^{2})\mapsto\lambda$, and $r\Lambda^{\eta_{1}-2}\mapsto r$. The dimensionless cutoff functions $f_{i}$, with $i\in\{p,p^{2},\omega,\omega^{2},r,\lambda,g\}$, arise from the integration over Matsubara frequencies and are normalized such that $f_{i}(1)=1$. For the sharp momentum cutoff used here, they take the values $f_{p}(x)=4(2+x)/[3(1+x)^{2}]$, $f_{p^{2}}(x)=f_{r}(x)=1/x$, $f_{\omega}(x)=4/(1+x)^{2}$, $f_{\omega^{2}}(x)=f_{\lambda}(x)=1/x^{3}$, and $f_{g}=2/[x(1+x)]$.
Figure 7: (a) Renormalization group flow in the plane spanned by the velocity ratio $v_{\mathrm{F}}/v_{\mathrm{B}}$ and the Yukawa coupling $g^{2}$, with the tuning parameter $r$ set to criticality, $r=r_{\star}$. Arrows indicate the flow toward the infrared. The stable fixed point at $v_{\mathrm{F}}/v_{\mathrm{B}}=1$ and $g^{2}=4\varepsilon/5$ belongs to the Gross-Neveu-Heisenberg (GNH) universality class. (b) Velocity ratio $v_{\mathrm{F}}/v_{\mathrm{B}}$ (blue) and effective dynamical critical exponent $z$ (red) as functions of the renormalization group scale $1/b$ for $v_{\mathrm{F}}/v_{\mathrm{B}}=1.35$ at the ultraviolet scale ($1/b=1$). In the infrared limit ($1/b\to 0$), $v_{\mathrm{F}}/v_{\mathrm{B}}$ and $z$ both flow to unity, reflecting the emergence of Lorentz invariance at criticality.
The flow equations exhibit a unique critical fixed point at $g^{2}_{\star}=\frac{4}{5}\varepsilon+\mathcal{O}(\varepsilon^{2})$ and $\lambda_{\star}=\frac{8}{55}\varepsilon+\mathcal{O}(\varepsilon^{2})$, illustrated in Fig. 7(a). Importantly, the fixed point is characterized by a single velocity,
$\frac{v_{\mathrm{F}}}{v_{\mathrm{B}}}$ $=1\,$, (37)
so that both fermionic and magnetic excitations propagate at the same speed at low energies. As a result, the system undergoes a continuous Dirac-semimetal-to-antiferromagnetic-insulator transition exhibiting emergent Lorentz invariance [81]. This is illustrated in Fig. 7(b), which shows the evolution of the velocity ratio $v_{\mathrm{F}}/v_{\mathrm{B}}$ (blue) and the effective dynamical exponent $z$ from Eq. (34) (red) as functions of the renormalization group scale $1/b$, with $1/b=1$ and $1/b\to 0$ corresponding to the ultraviolet and infrared limits, respectively. Here, we initialize the flow with a velocity ratio consistent with the moiré-scale estimate $v_{\mathrm{F}}/v_{\mathrm{B}}=1.35$, and choose the bare couplings $g^{2}=g_{\star}^{2}$ and $\lambda=0$ at the ultraviolet scale. Other choices of starting values produce qualitatively similar flows and the same infrared limit, provided that $r$ is tuned to criticality. Using Eqs. (34)–(36) together with the fixed-point values, we obtain the critical exponents that characterize the universal behavior near the quantum critical point as
$z$ $=1\,$, $\nu$ $=\frac{1}{2}+\frac{21\varepsilon}{55}+\mathcal{O}(\varepsilon^{2})\,$, (38)
$\eta_{\psi}$ $=\frac{3\varepsilon}{10}+\mathcal{O}(\varepsilon^{2})\,$, $\eta_{\phi}$ $=\eta_{1}=\frac{4\varepsilon}{5}+\mathcal{O}(\varepsilon^{2})\,.$ (39)
These values agree precisely with the established results for Gross-Neveu-Heisenberg universality to leading order in the $\varepsilon$ expansion about the upper critical dimension [3, 5]. This explicitly confirms that the semimetal-to-insulator transition in twisted double bilayer WSe2 belongs to the Gross-Neveu-Heisenberg universality class.
To go beyond the leading-order $\varepsilon$ expansion, we can therefore rely on established results for Gross-Neveu-Heisenberg criticality to predict the quantum critical behavior. The critical behavior of the Gross-Neveu-Heisenberg model has been studied extensively using a variety of analytical approaches, including the $\varepsilon$ expansion around the upper [3, 82] and lower [83] critical dimensions, the large-$N$ expansion [84], and the functional renormalization group [5, 85, 86]. Complementary numerical investigations have explored the transition using different lattice realizations [4, 87, 88, 89, 6, 90, 91, 92, 93, 94, 95, 96].
These results enable predictions of the universal power laws governing the thermodynamic, spectroscopic, and transport properties near the quantum critical point. In the antiferromagnetic phase, the staggered magnetization scales with twist angle or pressure as $\langle\vec{\varphi}\rangle\propto(\theta_{\mathrm{c}}-\theta)^{\beta}$ or $\langle\vec{\varphi}\rangle\propto(p-p_{\mathrm{c}})^{\beta}$, with nonuniversal prefactors but a universal order-parameter exponent $\beta\approx 1.21$. The electronic gap exhibits the scaling $\Delta\propto(\theta_{\mathrm{c}}-\theta)^{\nu z}$ or $\Delta\propto(p-p_{\mathrm{c}})^{\nu z}$, with dynamical exponent $z=1$ and correlation-length exponent $\nu\approx 1.20$. At the quantum critical point $\theta=\theta_{\mathrm{c}}$ or $p=p_{\mathrm{c}}$, the fermionic (magnetic) excitation spectrum will be characterized by a continuum of low-energy excitations near the $\boldsymbol{\kappa}$ ($\boldsymbol{\gamma}$) point of the moiré Brillouin zone, with lower bounds
$\varepsilon_{\mathbf{k},\text{F}}$ $\geq v_{\mathrm{F}}|\mathbf{k}-\boldsymbol{\kappa}|+\mathcal{O}((\mathbf{k}-\boldsymbol{\kappa})^{2})\,$, (40)
$\varepsilon_{\mathbf{k},\text{B}}$ $\geq v_{\mathrm{B}}|\mathbf{k}|+\mathcal{O}(\mathbf{k}^{2})\,.$ (41)
Importantly, in the low-energy limit the velocities $v_{\mathrm{F}}$ and $v_{\mathrm{B}}$ flow to the same value, reflecting the emergent Lorentz invariance at criticality. Furthermore, the dynamical spin structure factor at the quantum critical point scales as $\mathcal{S}(\mathbf{k},\omega)\propto 1/{(\omega^{2}-\mathbf{k}^{2})^{(2-\eta_{\varphi})/2}}$, with order-parameter anomalous dimension $\eta_{\varphi}\approx 1.01$. The quoted exponents are estimated from interpolations between expansions near the lower and upper critical dimensions [83], using standard hyperscaling relations [97].
@@ VIII.2 Strained samples
We emphasize that the O(3) model above is valid only at energies below the strain-induced single-particle gap $\Delta_{\epsilon}$. For weakly strained samples, $\Delta_{\epsilon}$ is small compared to the microscopic interaction scale $J$, which sets the order of the interaction-induced band gap $\Delta$ deep in the ordered phase. In the strong-coupling regime, this scale can be estimated as $J\sim 4\overline{t}^{2}/U_{\text{S}}$, where $\overline{t}$ is the $\mathbf{k}$-averaged bandwidth and $U_{\text{S}}$ is the effective interaction strength responsible for the antiferromagnetic order, see Appendix B.3. For twist angles near the critical value, $\theta\simeq\theta_{\mathrm{c}}=2.7^{\circ}$, we estimate $\overline{t}\simeq 4.8\,\text{meV}$ and $U_{\mathrm{S}}\simeq 4.6\,\text{meV}$ for $\epsilon_{\text{eff}}=110$, giving $J\simeq 20\,\text{meV}$ for the microscopic energy scale. At energies above $\Delta_{\epsilon}$ but below $J$, such systems remain governed by Gross-Neveu-Heisenberg universality, described by Eq. (30). This leads to a crossover within the quantum critical regime at finite temperatures $T$ above the quantum critical point at $\theta=\theta_{\mathrm{c}}$ or $p=p_{\mathrm{c}}$, as schematically shown in Fig. 8.